OpenStudy (nevillesnetwork):

Medal & Fan (plz help asap) A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item: C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (nevillesnetwork):

@StudyGurl14 @Luigi0210 @triciaal @Unavailabilityy Will you guys help me out?

OpenStudy (triciaal):

discount pay less additional more so add total is 100%

OpenStudy (nevillesnetwork):

I don't get what the 0.765x and 0.06 is in "C = 0.765x + 0.06(0.765x)"

OpenStudy (nes1234):

Plug in your equation, C. C = 0.765(100) + 0.06[0.765(100)] C = 76.5 + 0.06(76.5) C = 76.5 + 4.59 C = 81.09 After the discount, Chris will pay $81.09 for the coat.

OpenStudy (nes1234):

*cost

OpenStudy (nes1234):

C=0.765(100)+0.06[0.765(100)] C=$81.09 for 25% of discount the employee will now have to pay $20.273. ( The point of mentioning the discount 10% +15%= 25%)

OpenStudy (nevillesnetwork):

So the answer is $20.273. @nes1234

OpenStudy (triciaal):

I don't quite follow that equation

OpenStudy (triciaal):

do you have answer choices?

OpenStudy (nevillesnetwork):

No, its a "empty box" I have to type the answer in.

OpenStudy (triciaal):